OpenStudy (anonymous):

Find f(x) and g(x) so that the function can be described as y = f(g(x)). y = two divided by x squared + 9

OpenStudy (anonymous):

@Nnesha plssss

OpenStudy (anonymous):

@agent0smith

Nnesha (nnesha):

\[y=\frac{ 2 }{ x^2 }+9\] like this ?

OpenStudy (anonymous):

yes

Nnesha (nnesha):

f(g(x)) meaning substitute x for g(x) function

OpenStudy (anonymous):

yeah, and then you simplify?

Nnesha (nnesha):

right but this one is backwardz u have to find f(x) and g(x) from\[y=\frac{ 2 }{ x^2 }+9\]

OpenStudy (anonymous):

is f(x) = 2/x + 9 and g(x) = x^2

Nnesha (nnesha):

hmm

Nnesha (nnesha):

could be this \[\frac{ 2 }{ x^2 } +9 \] g(x) =x hmmm

OpenStudy (anonymous):

yeah do both work?

OpenStudy (anonymous):

i got it right, it was whtai wrote :)

OpenStudy (anonymous):

i have one more coudl you help wiht that?

Nnesha (nnesha):

hmm okay i'll try

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f(x) = Quantity x plus four divided by six and g(x) = 6x - 4

OpenStudy (anonymous):

f(x) = x+4 / 6 g(x) = 6x - 4

Nnesha (nnesha):

alright first try f(g(x))=x substitute x for g(x) function \[\huge\rm f(\color{ReD}{g(x}))=\frac{( \color{ReD}{6x-4} )+ 4}{ 6 }\] now solve right side is it equal to x ??

OpenStudy (anonymous):

yes

Nnesha (nnesha):

alright now 2nd one g(f(x) ) is this equal to x ?

OpenStudy (anonymous):

umm

OpenStudy (anonymous):

yes?

Nnesha (nnesha):

i don't know it yet

Nnesha (nnesha):

substitute x for f(x) function into f(x)

OpenStudy (anonymous):

ok

Nnesha (nnesha):

\[\huge\rm g(\color{ReD}{f(x)})=6\color{ReD}{x}-4\] replace x with (x+4)/6

OpenStudy (anonymous):

6(x+4)/6 -4

Nnesha (nnesha):

right \[6(\frac{x+4 }{ 6 })-4\] is it equal to x ?

OpenStudy (anonymous):

yers

Nnesha (nnesha):

both are equal that shows both are inverse of each other f(g(x)) = g(f(x))

OpenStudy (anonymous):

thanks so muhc!

Nnesha (nnesha):

np :=)